# Pcc-tuning: Breaking the Contrastive Learning Ceiling in Semantic Textual Similarity

### v1

@@ Abstract
Semantic Textual Similarity (STS) constitutes a critical research direction in computational linguistics and serves as a key indicator of the encoding capabilities of embedding models. Driven by advances in pre-trained language models and contrastive learning techniques, leading sentence representation methods can already achieved average Spearman’s correlation scores of approximately 86 across seven STS benchmarks in SentEval. However, further improvements have become increasingly marginal, with no existing method attaining an average score higher than 87 on these tasks. This paper conducts an in-depth analysis of this phenomenon and concludes that the upper limit for Spearman’s correlation scores using contrastive learning is 87.5. To transcend this ceiling, we propose an innovative approach termed Pcc-tuning, which employs Pearson’s correlation coefficient as a loss function to refine model performance beyond contrastive learning. Experimental results demonstrate that Pcc-tuning markedly surpasses previous state-of-the-art strategies, raising the Spearman’s correlation score to above 90. 111Our code and checkpoints are available at https://github.com/ZBWpro/Pcc-tuning.
@@ 1 Introduction
Sentence embeddings are vector encodings that encapsulate the semantic essence of original texts. Owing to their capacity to facilitate offline computation as well as their pivotal role in realizing retrieval-augmented generation (Zhao et al., 2024), research in this area has garnered considerable attention from numerous institutions and scholars in recent years.
The quality of sentence embeddings is typically assessed via the SentEval (Conneau and Kiela, 2018) toolkit, which measures models based on their average Spearman correlation across seven STS benchmarks. With the continuous advancement of pre-trained language models (PLMs), contrastive learning, and prompt engineering, cutting-edge work in this field has elevated the scores on the leaderboard from an initial 60 (Pennington et al., 2014) to about 86 (Jiang et al., 2023b). As a result, the "PLM + contrastive learning" framework has become the mainstream paradigm in sentence representation research.
However, as illustrated in Table 1, models’ performance on standard STS tasks in SentEval appears to have hit a significant bottleneck. Whether utilizing classical discriminative PLMs such as BERT (Devlin et al., 2019) or emerging generative PLMs like LLaMA2 (Touvron et al., 2023b) and Mistral (Jiang et al., 2023a), contemporary state-of-the-art (SOTA) strategies are unable to achieve Spearman’s correlation scores higher than 87. Moreover, despite variations in training datasets, contrastive learning loss functions, and model architectures, the final scores are generally similar if the same type of PLM is selected.
@@
In this regard, Li and Li (2023b) posit that PLMs may have reached their performance limits in STS tasks. However, this paper will demonstrate through rigorous mathematical derivation that the core factor causing this performance ceiling is not the inadequacy of PLMs, but the inherent flaws in contrastive learning loss functions. Specifically, contrastive learning only distinguishes between two categories: similar and dissimilar, in determining the semantic relationships between text pairs. This binary classification strategy restricts its maximum achievable Spearman’s correlation score to 87.5, even under optimal conditions.
Following this proof, we introduce Pcc-tuning, a novel approach that employs a two-stage training process. This method enhances models’ semantic discrimination capabilities by utilizing a small amount of fine-grained annotated data post contrastive learning. With the same 7B-scale generative PLM, Pcc-tuning can achieve an average Spearman’s correlation score exceeding 90 on the aforementioned seven STS tasks, significantly surpassing previous best results.
@@
By analyzing the theoretical limits of binary classifiers in STS tasks, we demonstrate that the upper bound of Spearman’s correlation scores using contrastive learning methods is 87.5. This finding effectively explains the performance bottlenecks encountered by prior sentence representation strategies.
Building upon this, we propose Pcc-tuning, a method capable of taking full advantage of fine-grained labeled data with Pearson correlation as its loss function. After fine-tuning PLMs with contrastive learning, we only need to introduce annotated text pairs amounting to 3.7% of the original training set to bring notable performance improvements.
We extensively validate the effectiveness of Pcc-tuning across internationally recognized STS benchmarks and seven transfer tasks. Experimental results show that Pcc-tuning significantly outperforms previous SOTA methods across different PLMs and prompts.
@@ 2.1 Contrastive Learning and Binary Classifiers
Currently, leading approaches for sentence representation predominantly center around contrastive learning, with InfoNCE Loss (Oord et al., 2018) being the most commonly adopted loss function. Given an input text , InfoNCE Loss computes the similarity between this sample and its positive example in the numerator, and contrasts it with the similarity calculations between and other texts within the same batch in the denominator. This formulation aims to bring similar instances closer while pushing dissimilar ones apart. The mathematical expression for InfoNCE Loss is presented in Equation 1, where denotes the encoding method, represents the batch size, and signifies a temperature hyperparameter.
Equation 1 reveals that contrastive learning loss functions, exemplified by InfoNCE Loss, essentially classify sentence pairs into two distinct classes: similar and dissimilar. However, no further distinctions are made within these two categories. In other words, as long as is semantically different from or , InfoNCE Loss treats both and as negative sample pairs. As for which of and exhibits a lower degree of similarity, contrastive learning neither concerns itself with this information nor can it readily leverage such details. Indeed, for the majority of embedding models, their training sets are specially adjusted to provide coarse-grained categorical annotations, so as to better align with the contrastive learning framework (Gao et al., 2021).
@@ 2.3 The Spearman Correlation Upper Limit of Contrastive Learning Methods
As discussed in Section 2.1, contrastive learning differentiates texts based on binary semantic relations: similar and dissimilar. Thus, its effectiveness parallels that of a binary classifier. This section derives the optimal Spearman correlation achievable by a binary classifier in STS tasks, thereby elucidating the performance upper bound of contrastive learning methods.
Given a collection of text pairs comprising samples, we initially arrange the elements of in descending order according to manually annotated semantic similarity, yielding the sorted set . Assume that . Then, for any binary classifier, its performance reaches the optimum only when it categorizes the first sample pairs of as positive examples and the remaining sample pairs as negatives. Otherwise, it indicates at least one misclassification.
Since this binary classifier is solely responsible for constructing an optimal classification boundary between the two categories of similarity and dissimilarity (i.e., distinguishing only whether two texts are semantically akin), its predicted score for the first samples is consistently identical (assumed to be 1), and likewise for the last samples (assumed to be 0). By the definition of Spearman’s correlation coefficient, the difference in rankings between predictions and true values, , alongside , can be represented as:
These equations showcase that can be viewed as a function of . Upon rearranging, we derive: (more details can be found in Appendix C.)
In Equation 5, remains constant, thus is contingent on . When , i.e., when the model deems the first of sample pairs as positives and the latter as negatives, attains its minimum. Therefore, the minimum value of is:
@@
Apart from the original InfoNCE Loss, an extended contrastive learning loss function tailored for NLI datasets (Bowman et al., 2015; Williams et al., 2018), as shown in Formula 8, is frequently utilized in sentence representation research (Gao et al., 2021; Zhang et al., 2023). The incorporation of hard negative example in the denominator, equivalent to enlarging the batch size, does not affect the correctness of our derivation.
It should be noted that the above conclusion has been validated through numerous experiments. To date, embedding derivation schemes based on contrastive learning have not achieved a Spearman’s correlation score above 87. This theoretical analysis provides clear guidance for this empirical observation.
@@ 3 Proposed Method
This section introduces Pcc-tuning, an innovative solution for STS tasks. Pcc-tuning employs a two-stage training pipeline and is designed to surpass the 87.5 performance upper bound of contrastive learning methods.
@@
The anisotropy of PLMs’ semantic space (Ethayarajh, 2019) presents a longstanding challenge in sentence representation research. Contrastive learning has proven effective in stabilizing embedding distances among semantically similar texts while ensuring a more uniform distribution of vector encodings (Gao et al., 2021), thereby markedly enhancing the semantic space properties of PLMs. Consequently, leveraging contrastive learning to refine the initial state of pre-trained models has emerged as a prevalent strategy within the NLP community (Wang et al., 2022; Li et al., 2023).
@@
Following this established practice, we initially conduct supervised fine-tuning of the PLM using the NLI dataset constructed by SimCSE (Gao et al., 2021). This dataset comprises 275,602 text pairs in triplet form, providing a robust source of coarse-grained labeled information for the model. Our implementation in the first stage closely aligns with that of PromptEOL (Jiang et al., 2023b), where we load the original PLM checkpoint and fine-tune the model with the extended InfoNCE Loss depicted in Equation 8, combined with QLoRA (Dettmers et al., 2024). A unique feature of our methodology is the adoption of the PromptSTH template proposed by Zhang et al. (2024): "This sentence : ‘[X]’ means something", which encapsulates the input sentence [X] and extracts the encoding of the final token as the sentence embedding. Later sections will examine the performance of Pcc-tuning under various prompts.
@@
After the contrastive learning phase, the semantic space of the PLM will be adjusted to a superior encoding state, capable of generating high-quality embeddings. However, the inability of InfoNCE Loss to harness fine-grained annotation information leads to a pronounced performance bottleneck for contrastive learning methods in STS tasks. To mitigate this issue, a finer distinction is required within the two categories of similarity and dissimilarity, along with introducing the ordinal relationships of text pairs in terms of semantic similarity.
@@
The optimal strategy is to incorporate fine-grained annotated data in the second stage and guide the model’s training process via Spearman’s correlation coefficient. This ensures maximum consistency between the model’s behavior during training and testing phases. However, as Spearman correlation is non-differentiable and thus incompatible with backpropagation, we opt for Pearson’s correlation coefficient to update model parameters, which also serves as the inspiration for the name Pcc-tuning. Pearson correlation and our loss function in the second stage are shown in Equation 9, where represents the cosine similarity between model-derived embeddings, and denotes the human-annotated scores for the text pairs.
@@
Employing Pearson coefficient as the loss function enables effective utilization of fine-grained annotation information and supports diverse combinations with a small volume of data. For instance, the tuning dataset in our second stage consists of the training sets of STS-B (Cer et al., 2017) and SICK-R (Marelli et al., 2014), which together contain 10,249 text pairs. This number merely represents 3.7% of the first-stage training dataset, yet their combination varieties reach up to . Therefore, even with multiple epochs of training, the similarity ranking of samples in each batch is unlikely to repeat.
Figure 2 provides a detailed illustration of Pcc-tuning’s training process. In the first stage, we fine-tune the model using contrastive learning and the NLI dataset. In the second stage, we introduce a small amount of fine-grained annotated data and load the checkpoint from the first phase to further update the model parameters via Pearson’s correlation coefficient.
@@ 4 Experiments
This section presents the experimental results of Pcc-tuning. Initially, in subsection 4.1, we elaborate on our experimental setup, including evaluation methods, datasets, and the selection of baselines. Subsequently, in subsection 4.2, we compare the performance of Pcc-tuning with contemporary SOTA text representation strategies on internationally recognized STS benchmarks. Finally, in subsection 4.3, we validate the efficacy of Pcc-tuning under diverse prompts.
@@ 4.1 Implementation Details
In line with prior studies (Gao et al., 2021; Jiang et al., 2022, 2023b), we utilize the SentEval (Conneau and Kiela, 2018) toolkit to assess our model across seven STS tasks, with Spearman’s correlation coefficient as the core metric.
As outlined in Section 3, Pcc-tuning incorporates a two-stage training pipeline. The respective training sets originate from the NLI dataset organized by SimCSE (Gao et al., 2021), containing 275,602 text pairs, and a mixed dataset composed of the training sets from STS-B and SICK-R, totaling 10,249 text pairs. In all experiments, only during the testing phase can models access data from the evaluation benchmarks. It is noteworthy that although Pcc-tuning requires specific corpora at both stages, the total data volume employed is only 285,851 entries. In contrast, the publicly available training data for the current SOTA method, DeeLM (Li and Li, 2023b), includes 480,862 triplet text pairs, with additional data remaining inaccessible.
Our experiments are conducted using several widely adopted 7B-scale generative PLMs: OPT6.7b (Zhang et al., 2022), LLaMA7b (Touvron et al., 2023a), LLaMA27b, and Mistral7b. To clearly demonstrate the superiority of Pcc-tuning, we primarily compare it against current SOTA strategies. Specifically, among our selected baselines, PromptEOL (Jiang et al., 2023b), PromptSTH (Zhang et al., 2024), AngIE (Li and Li, 2023a), and DeeLM (Li and Li, 2023b) are leading generative PLM sentence representation methods, which significantly outperform BERT-based approaches on STS benchmarks. Meanwhile, openai-ada-002, jina-base-v2 (Günther et al., 2023), and nomic-embed-v1 (Nussbaum et al., 2024) represent the most advanced contrastive learning pre-trained models at present.
4.2 Main Results
Table 2 summarizes the results of the above experiments. Under all tested PLMs, Pcc-tuning consistently transcends the 87.5 Spearman correlation upper bound of contrastive learning methods, achieving an impressive average score of approximately 90. Notably, when employing Mistral7b as the backbone, Pcc-tuning attains a Spearman’s correlation score of 90.61, substantially surpassing the previous record of 86.01 set by DeeLM. Moreover, Pcc-tuning excels beyond prior SOTA methods in each of the seven STS tasks aggregated within SentEval, manifestly affirming its efficacy. These outcomes collectively underscore the crucial role of modeling fine-grained annotated information in STS tasks.
Furthermore, since Pcc-tuning’s first-stage implementation mirrors that of PromptSTH, the comparison between Pcc-tuning and PromptSTH in Table 2 also functions as an ablation study. It reveals that, constrained by the coarse granularity of contrastive learning, whether adopting the earlier released OPT model or the newly open-sourced Mistral model, the Spearman’s correlation scores for PromptSTH are confined between 85.3 to 85.7, showing limited progress. In contrast, Pcc-tuning provides improvements of about 5 percentage points, reaffirming the mathematical derivations discussed in Section 2.
@@ 4.2 Main Results
| PromptSTH | 78.48 | 90.09 | 85.10 | 88.71 | 85.93 | 88.51 | 80.95 | 85.40 |
| Pcc-tuning | 84.40 | 94.40 | 93.15 | 93.49 | 88.62 | 90.86 | 87.08 | 90.29 |
@@
| PromptSTH | 79.19 | 89.70 | 85.07 | 88.88 | 86.65 | 88.20 | 81.95 | 85.66 |
| Pcc-tuning | 85.77 | 93.79 | 93.78 | 94.02 | 89.07 | 90.73 | 87.14 | 90.61 |

### v2

@@ Abstract
Semantic Textual Similarity (STS) constitutes a critical research direction in computational linguistics and serves as a key indicator of the encoding capabilities of embedding models. Driven by advances in pre-trained language models and contrastive learning, leading sentence representation methods have reached an average Spearman’s correlation score of approximately 86 across seven STS benchmarks in SentEval. However, further progress has become increasingly marginal, with no existing method attaining an average score higher than 86.5 on these tasks. This paper conducts an in-depth analysis of this phenomenon and concludes that the upper limit for Spearman’s correlation scores under contrastive learning is 87.5. To transcend this ceiling, we propose an innovative approach termed Pcc-tuning, which employs Pearson’s correlation coefficient as a loss function to refine model performance beyond contrastive learning. Experimental results demonstrate that Pcc-tuning can markedly surpass previous state-of-the-art strategies with only a minimal amount of fine-grained annotated samples. 111Our code and checkpoints are available at https://github.com/ZBWpro/Pcc-tuning.
@@ 1 Introduction
Sentence embeddings refer to vector encodings that encapsulate the semantic essence of original texts. Owing to their capacity to facilitate offline computation as well as their pivotal role in realizing retrieval-augmented generation (Zhao et al., 2024), research in this area has garnered considerable attention from numerous institutions and scholars in recent years.
The quality of sentence embeddings is typically assessed via the SentEval (Conneau and Kiela, 2018) toolkit, which measures models based on their average Spearman correlation across seven STS benchmarks. With the continuous advancement of pre-trained language models (PLMs), contrastive learning, and prompt engineering, cutting-edge works in this field have progressively elevated leaderboard scores from an initial 60 (Pennington et al., 2014) to around 86 (Jiang et al., 2023b). As a result, the "PLM + contrastive learning" framework has become the mainstream paradigm in sentence representation research.
However, as illustrated in Table 1, models’ performance on standard STS tasks in SentEval appears to have hit a significant bottleneck. Whether utilizing classical discriminative PLMs like BERT (Devlin et al., 2019) or emerging generative PLMs such as LLaMA2 (Touvron et al., 2023b) and Mistral (Jiang et al., 2023a), contemporary state-of-the-art (SOTA) strategies are unable to achieve Spearman’s correlation scores exceeding 86.5. Moreover, despite variations in training datasets, contrastive learning objectives, and model architectures, the final performance are generally similar if the same type of PLM is selected.
@@
In this regard, DeeLM (Li and Li, 2023b) posits that PLMs may have reached their performance limits on STS tasks. However, this paper will demonstrate through rigorous mathematical derivation that the core factor causing this performance ceiling is not the inadequacy of PLMs, but inherent flaws in contrastive learning loss functions. Specifically, contrastive learning only distinguishes between two categories: similar and dissimilar, in determining the semantic relationships between text pairs. This binary classification strategy restricts its maximum achievable Spearman’s correlation score to 87.5, even under optimal conditions.
Following this proof, we introduce Pcc-tuning, a novel approach that employs a two-stage training process. This method enhances models’ semantic discrimination capabilities by leveraging a small amount of fine-grained annotated data post contrastive learning. With the same 7B-scale generative PLMs, Pcc-tuning can substantially surpass previous best results on the seven aforementioned STS tasks and break through the performance ceiling of 87.5.
@@
By analyzing the theoretical limits of binary classifiers in STS tasks, we prove that the upper bound for Spearman’s correlation scores using contrastive learning methods is 87.5. This finding effectively explains the performance plateau encountered by prior sentence representation research.
Building upon this, we propose Pcc-tuning, a method capable of taking full advantage of fine-grained labeled data with Pearson correlation as its loss function. After fine-tuning PLMs through contrastive learning, we only need to introduce annotated text pairs amounting to 1.96% of the original training set to bring notable performance improvements.
We extensively validate the effectiveness of Pcc-tuning across internationally recognized STS benchmarks and multiple transfer tasks. Experimental results show that Pcc-tuning consistently outperforms existing SOTA methods across different PLMs, prompts and hyperparameter settings.
@@ 2.1 Contrastive Learning and Binary Classifiers
Currently, leading approaches for sentence representation predominantly center around contrastive learning, with InfoNCE Loss (Oord et al., 2018) being the most commonly adopted loss function. Given an input text , InfoNCE Loss computes the similarity between this sample and its positive example in the numerator, contrasting it with the similarity calculations between and other texts within the same batch in the denominator. This formulation aims to bring similar instances closer while pushing dissimilar ones apart. The mathematical expression for InfoNCE Loss is presented in Equation 1, where denotes the encoding method, represents the batch size, and signifies a temperature hyperparameter.
Equation 1 indicates that contrastive learning loss functions, exemplified by InfoNCE Loss, essentially classify sentence pairs into two distinct classes: similar and dissimilar. However, no further distinctions are made within these two categories. In other words, as long as is semantically different from or , InfoNCE Loss treats both and as negative sample pairs. As for which of and exhibits a lower degree of similarity, contrastive learning neither concerns itself with this information nor can it readily leverage such details. Indeed, for the majority of embedding models, their training sets are specially adjusted to provide coarse-grained categorical annotations, so as to better align with the contrastive learning framework (Gao et al., 2021; Jiang et al., 2022, 2023b).
@@
One potential concern with this analogy is that contrastive learning models compute cosine similarities between embeddings during the testing phase, resulting in continuous predicted values. Nevertheless, since the model does not differentiate between internal discrepancies within the positive and negative classes during training, it cannot be expected to possess the capability to discern fine-grained semantic similarities. As training data continues to flow in, InfoNCE Loss gradually guides the model toward the characteristics of an ideal binary classifier. However, constrained by the expressive power of neural networks, as well as the scale and quality of the training data, relying solely on contrastive learning is insufficient to replicate the performance of a binary classifier that is free from inversely ordered pairs. Therefore, it is reasonable to consider the optimal binary classifier as the ultimate state of contrastive learning models.
@@ 2.3 The Spearman Correlation Upper Limit of Contrastive Learning Methods
As discussed in section 2.1, contrastive learning distinguishes texts based on coarse-grained semantic relations, categorizing them as either similar or dissimilar. Thus, its effectiveness parallels that of a binary classifier. This section derives the optimal Spearman correlation achievable by a binary classifier in STS tasks, thereby elucidating the performance upper bound of contrastive learning methods.
Given a collection of text pairs consisting of samples, we initially arrange the elements of in descending order according to manually annotated semantic similarity, yielding the sorted set . Assume that . Then, for any binary classifier, its performance reaches the optimum only when it categorizes the first sample pairs of as positive examples and the remaining sample pairs as negatives. Otherwise, it indicates at least one misclassification.
Since this binary classifier is solely responsible for constructing an optimal classification boundary between the two categories of similarity and dissimilarity (i.e., determining only whether two texts are semantically akin), its predicted scores for the first samples are consistently identical (assumed to be 1), and likewise for the last samples (assumed to be 0). By the definition of Spearman’s correlation coefficient, the difference in rankings between predictions and true values, , alongside , can be represented as:
These equations showcase that can be viewed as a function of . Upon rearranging, we derive: (with further details provided in Appendix A.)
In Equation 4, remains constant, so depends on . When , i.e., when the model deems the first of sample pairs as positives and the remaining as negatives, attains its minimum. Therefore, the minimum value of is:
@@
Apart from the original InfoNCE Loss, an extended contrastive learning loss function tailored for NLI datasets (Bowman et al., 2015; Williams et al., 2018), as shown in Formula 7, is frequently utilized in sentence representation research (Gao et al., 2021; Zhang et al., 2024a). The incorporation of hard negative example in the denominator, which is equivalent to enlarging the batch size, does not affect the correctness of our derivation.
It should be noted that the above conclusion has been validated through numerous experiments. To date, embedding derivation schemes based on contrastive learning have not achieved a Spearman’s correlation score above 86.5. This theoretical analysis provides a clear explanation for these empirical observations.
@@ 3 Proposed Method
This section introduces Pcc-tuning, an innovative strategy for addressing STS tasks. Pcc-tuning employs a two-stage training pipeline and is designed to break the 87.5 performance ceiling in contrastive learning methods.
@@
The anisotropy of PLMs’ semantic space (Ethayarajh, 2019) is a longstanding challenge in sentence representation research. Contrastive learning has proven effective in stabilizing embedding distances among semantically similar texts while promoting a more uniform distribution of overall vector encodings (Gao et al., 2021), thus markedly enhancing the semantic properties of PLMs. Therefore, leveraging contrastive learning to refine the initial state of pre-trained models has emerged as a prevalent approach within the NLP community (Wang et al., 2022; Li et al., 2023; Muennighoff et al., 2024).
@@
Following this well-established practice, we initially conduct supervised fine-tuning of the PLM using the NLI dataset constructed by SimCSE (Gao et al., 2021). This dataset comprises 275,601 triplet-form text pairs, providing a robust source of coarse-grained labeled information for the model. Our implementation in the first stage closely mirrors that of PromptEOL (Jiang et al., 2023b) for comparison purposes, where we load the original PLM checkpoint and fine-tune the model with the extended InfoNCE Loss depicted in Equation 7, combined with QLoRA (Dettmers et al., 2024). A distinctive feature of our methodology is the adoption of the PromptSUM template proposed by Zhang et al. (2024b): "This sentence : ‘[X]’ can be summarized as", which encapsulates the input sentence [X] and extracts the encoding of the final token as the sentence embedding. Later sections will examine Pcc-tuning’s performance under various prompts.
@@
After the contrastive learning phase, the PLM will be adjusted to a superior encoding state, capable of producing high-quality embeddings. However, the neural network trained at this stage remains insufficient as the final solution for STS tasks. This is due to two primary reasons: (1) The contrastive learning objective does not fully align with the evaluation metrics of STS tasks. While a decrease in InfoNCE Loss reflects a better clustering effect of sentence vectors in semantic space, this does not necessarily translate into an improvement in Spearman correlation. The latter essentially measures the consistency of the model’s scoring with human ratings in terms of monotonicity. (2) As discussed in section 2, contrastive learning loss functions fail to harness fine-grained annotation information, leading to a pronounced performance bottleneck. Consequently, the benefits of further optimizing binary classification performance diminish with successive iterations. To mitigate these issues, a finer distinction is required within the two categories of similarity and dissimilarity, along with introducing ordinal relationships of text pairs based on semantic similarity.
@@
The optimal strategy is to incorporate fine-grained annotated data in the second stage and guide the model’s training process via Spearman’s correlation coefficient. This ensures maximum consistency between the model’s behavior during training and testing. However, since Spearman correlation is non-differentiable and thus incompatible with backpropagation, we opt for Pearson’s correlation coefficient to update model parameters, which also serves as the inspiration for the name Pcc-tuning. Pearson correlation and our loss function for the second stage are shown in Equation 8, where represents the cosine similarity between model-derived embeddings, and denotes the human-annotated scores for the text pairs.
@@
Employing Pearson’s correlation coefficient as the loss function enables effective utilization of fine-grained sample scores and supports diverse combinations even with relatively small data quantities. For example, our tuning dataset in the second stage is composed of filtered training sets from STS-B (Cer et al., 2017) and SICK-R (Marelli et al., 2014), which together contain 5,398 text pairs. This number merely constitutes 1.96% of the size of the NLI dataset adopted in the first stage, yet the potential combination varieties reach up to (where represents the batch size). As a result, even after multiple epochs of training, the similarity rankings of samples in each batch are unlikely to repeat, thereby continuously providing the model with meaningful gradient information.
The dataset is filtered because we discovered that some sentence pairs in the STS-B and SICK-R training sets overlap with the test sets of the seven STS benchmarks in SentEval. Additionally, there is even overlap between the train and test sets within SICK-R itself. To prevent information leakage, we implemented a stringent filtering mechanism to ensure that the model does not encounter any test set text pairs during parameter updates. More details about this filtering process can be found in Appendix B.
Figure 2 presents an overview of Pcc-tuning’s training pipeline. In the first stage, the model is fine-tuned using contrastive learning on the NLI corpus. In the second stage, we introduce a small amount of fine-grained annotated data and load the checkpoint from the first phase to further update the model parameters via Pearson’s correlation coefficient. This two-stage fine-tuning strategy effectively prevents overfitting. Although the filtered STS-B and SICK-R training sets provide only 5,398 fine-grained labeled instances, the 275,601 text pairs used in the first stage establish a solid initial state for the model, thereby maximizing its generalization capacity. Moreover, the varying scoring scales of STS-B and SICK-R also introduce a degree of noise into the Pcc-tuning training process , which contributes to the model’s robustness. We provide detailed results of our method’s zero-shot performance on several downstream tasks in Appendix C to further demonstrate its transferability.
@@ 4 Experiments
This section presents the experimental results of Pcc-tuning. Initially, in subsection 4.1, we elaborate on our experimental setup, including evaluation methods, training data size, and the selection of baselines. Subsequently, in subsection 4.2, we compare the performance of Pcc-tuning with current SOTA text representation strategies across internationally recognized STS benchmarks. Following this, in subsection 4.3, we conduct targeted experiments to demonstrate that Pcc-tuning surpasses continuous contrastive learning. Finally, in subsection 4.4, we examine the efficacy of Pcc-tuning under different prompts.
@@ 4.1 Implementation Details
In line with prior studies (Gao et al., 2021; Jiang et al., 2022, 2023b; Chen et al., 2023), we utilize the SentEval (Conneau and Kiela, 2018) toolkit to assess our model on seven STS tasks, with Spearman’s correlation coefficient as the core metric. In all experiments, models are permitted access to text pairs from the evaluation benchmarks only during the testing phase.
It is noteworthy that although Pcc-tuning requires specific corpora at both stages of training, the total data volume employed is only 280,999 entries. In contrast, the publicly available training data for the contemporary SOTA method, DeeLM (Li and Li, 2023b), includes 480,862 triplet text pairs, with additional datasets remaining inaccessible.
@@
| PromptSUM | 78.84 | 90.03 | 85.06 | 88.80 | 85.66 | 88.29 | 81.58 | 85.47 |
| Pcc-tuning (SUM) | 81.00 | 90.66 | 86.09 | 90.42 | 86.21 | 89.83 | 87.23 | 87.35 |
| PromptEOL | 79.00 | 89.80 | 85.10 | 88.86 | 86.03 | 88.48 | 81.06 | 85.48 |
| Pcc-tuning (EOL) | 81.41 | 91.15 | 86.62 | 90.69 | 86.99 | 89.97 | 86.85 | 87.67 |
@@
| PromptSUM | 79.76 | 89.69 | 85.33 | 89.30 | 86.62 | 88.27 | 81.81 | 85.83 |
| Pcc-tuning | 82.04 | 90.84 | 86.79 | 91.10 | 87.18 | 90.05 | 87.02 | 87.86 |
Our experiments are conducted based on several widely adopted 7B-scale generative PLMs: OPT6.7b (Zhang et al., 2022), LLaMA7b (Touvron et al., 2023a), LLaMA27b, and Mistral7b. To clearly demonstrate the superiority of Pcc-tuning, we primarily compare it against current SOTA strategies. Specifically, among our selected baselines, PromptEOL (Jiang et al., 2023b), PromptSUM (Zhang et al., 2024b), AngIE (Li and Li, 2023a), and DeeLM (Li and Li, 2023b) are leading generative PLM-based sentence representation methods, which significantly outperform BERT-based approaches on STS benchmarks. Meanwhile, openai-ada-002, jina-base-v2 (Günther et al., 2023), and nomic-embed-v1 (Nussbaum et al., 2024) represent the most advanced contrastive learning pre-trained models at present.
4.2 Main Results
Table 2 summarizes the performance of various methods on the seven STS tasks collected in SentEval. Under all tested PLMs, Pcc-tuning consistently surpasses previous SOTA strategies, either approaching or exceeding the Spearman correlation upper limit of 87.5 for contrastive learning methods. Notably, when Mistral7b is selected as the backbone, Pcc-tuning attains a Spearman correlation of 87.86, which is 2.15% higher than the leaderboard record of 86.01 set by DeeLM, despite DeeLM using a much larger training corpus. Moreover, considering that Pcc-tuning delivers the best performance across all seven STS tasks, its effectiveness is self-evident. These outcomes collectively underscore the crucial role of modeling fine-grained annotated information in STS tasks.
Furthermore, since Pcc-tuning’s first-stage is implemented identically to PromptSUM, the comparison between Pcc-tuning and PromptSUM in Table 2 also functions as an ablation study. It reveals that, constrained by the coarse granularity of contrastive learning, whether adopting the earlier released OPT model or the newly open-sourced Mistral model, Spearman’s correlation scores for PromptSUM are confined around 85.5, showing limited progress. In contrast, Pcc-tuning provides an improvement of approximately 2 percentage points, reaffirming the mathematical derivations discussed in section 2.